ABC Audio sells headphones and would like to earn after-tax profits of S480 every week. Each set of headphones incurs variable costs of $6 and sells for $12. Rent and other fixed costs are $300 per week; the income tax rate is 20%. How many headphones must ABC sell per week to meet its profit goal? ales headphone sets/week
media%2F902%2F902ec799-bd9f-489f-ad71-41
Answer #1

Answer:-1)-Selling price per unit =$12

Variable cost per unit =$6

Contribution margin per unit= Selling price per unit - Variable cost per unit

=$12 per unit -$6 per unit =$6 per unit

ABC must sell to achieve target profit =(Fixed costs+ Target profit)/Contribution margin per unit

=($480+$300)/$6 per unit =130 headphones

2)-MOS sales in units = Total sales in units - BEP sales in units

=260000 units – 156000 units

=104000 units

MOS in sales dollars =104000 units*$3.50 per unit =$364000
Answer #2
300 + 480:(1-0.20)/6-12 =300 + 480: 0.80/6 =300 + 600/6 =150 headphone sets per week
source: Cost Management textbook
answered by: Laura